An empty swimming pool with a square floor dimension has a volume of 3y2 m3. Given the weight of air in the swimming pool is about 500 N and the air density is 1.29 kg/m3.

Find the length of the swimming pool’s floor, y.


1
Expert's answer
2021-09-25T13:09:13-0400

"P=mg=\rho Vg=\rho\cdot3y^2\cdot g \to y=\sqrt{P/(3\rho g)}="


"=\sqrt{500/(3\cdot1.29\cdot9.81)}=3.63\ (m)"